If number of molecules of H2 are double than that of O2 , the ratio of kinetic energy of hydrogen and that of oxygen at 300K is 
 
 

Grade:12th pass

1 Answers

Rituraj Tiwari
3 years ago
The average kinetic energy translation per molecule of all gases is same for all gases at same temperature and is given by,
K.E. = 3/2 kT

- Hence for hydrogen molecules
K.E.(H2) = 3/2 kT
K.E.(O2) = 1.5×1.380×300 = 621J

- Also for oxygen molecules
K.E.(O2) = 3/2 kT
K.E.(O2) = 1.5×1.380×300 = 621J

K.E.(H2)/K.E.(O2) = 621/621 = 1
- Therefore, Ratio of hydrogen to oxygen is 1.